10/15/99 Evening Shift

Peter Shanahan (shanahan@fnal.gov)
Fri, 15 Oct 1999 23:58:01 -0500

Shift Summary for eve shift, 15/Oct/99

On Shift: M. Timko, P. Shanahan

Hours of beam available at all intensities: 8.15
Hours of beam available at nominal intensity: 8.15
Hours of E832 data (special runs not included): 8.00
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:
Beam - a few aborts
DAQ - change of tapes, a few spills to crate 53

========================================================================
Runs ended this shift:
Run 14846 - 3.2M events - BEAM_799

Tapes brought to FCC this shift:

UPA079,UPA080,UPA081
UPB079,UPB080,UPB081
UPC079,UPC080,UPC081
UPS027

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Very smooth data taking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14846 16:00 Patty and Jeff Kalenbach reboot xktev28c. Ed
restarts serious alarm application.

14846 16:46 Tapes full. Change tapes. Serious alarm application
goes away all by itself.

14847 16:57 During init, CsI 1548 echoes back 0xa585, instead of
0xb585. Everything is OK on new pipe init w/DBC.

14847 1 17:01 Start new run. Alarm widget starts and survives.

17:15 DN puts new Ar-Ethane bottles online

14847 45 18:00 Crate 53 hangs trigger. Kfrend twice.

14847 21:45 Bug MCR about 8.5E12 for last ~1/2hour. Partly
Booster, partly MI. It starts to improve.

14847 247 22:30 Crate 53 hangs trigger. Kfrend.

14847 23:25 mount ONSPLIT tape UPF021 on ktev4

14847 299 23:42 xktev10c was white (instead of pink) for 4 seconds.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Note: Many of the histogram references need work.

CSI - timing was goofy in reference (or made with some strange trigger
set?). RC histogram somehow ended up in the main page of CsI plots.

DC - some plots in reference have frowny faces.

SCIA - hitmap looks very sparse in reference.

Please take a look at your favorite reference section.